# Achieving Near-Optimal Regret for Bandit Algorithms with Uniform Last-Iterate Guarantee

### v1

@@ 4.1 Main Algorithm and Main Results
Theorem 4.2 also shows that, for each phase, the spanner can be constructed by calling the oracle for only polynomial times. Compared to the computational efficiency of the algorithm in (Awerbuch and Kleinberg, 2008), the efficiency of our algorithm is only $d$ times larger than theirs.
4.2 Key Technique: Adaptive Barycentric Spanner
Algorithm 4 Adaptive barycentric spanner
if $|{\mathcal{B}}_{m-1}|<d$ then
The primary reason that the non-adaptive barycentric spanner technique (Awerbuch and Kleinberg, 2008) fails to work for the elimination approach is that it requires the knowledge of the space that the active arm set ${\mathcal{A}}_{m}$ spans. However, acquiring such knowledge is often difficult, because set ${\mathcal{A}}_{m}$, which may span a proper linear subspace of $\mathbb{R}^{d}$, varies for different phases and the number of active arms could be infinite. Our algorithm shown in Algorithm 4 can identify a barycentric spanner for active arm set ${\mathcal{A}}_{m}$ adaptively for each phase $m$, even if they do not span $\mathbb{R}^{d}$.
Algorithm 4 initializes a matrix $A=[a_{1},\ldots,a_{d}]$ whose column vectors are linearly independent, and $a_{i}=e_{i}/\sqrt{T_{m}}$. At the beginning, the algorithm solves (7) to find an empirical best arm $a^{\star}_{m}$ (line 4-4). Then, in two for-loops (line 4-4 and line 4-4), the algorithm tries to replace each column vector of $A$ with an active arm $a\in{\mathcal{A}}_{m}$ by invoking the subroutine LI-Argmax in Algorithm 5, while keeping column vectors of $A$ linearly independent. For each column $i$, LI-Argmax attempts to find $\max_{a\in{\mathcal{A}}_{m}}|\det(a,A_{-i})|$ [^5] It is equivalent to $\max_{a\in{\mathcal{A}}_{m}}|\left\langle{u,a}\right\rangle|$ where $u$ can be found by rank-one update. See (Zhu et al., 2022) for more details. by solving (5) with $a^{\star}_{m}$. If $\max_{a\in{\mathcal{A}}_{m}}|\det(a,A_{-i})|=0$, then, the algorithm returns Null, since there does not exist an active arm that can maintain the linear independence. In this case, the algorithm keeps $e_{i}/\sqrt{T_{m}}$ in the $i$-th column of $A$. Otherwise, $a_{i}$ will be updated to $s_{i}\in\operatorname*{\arg\!\max}_{a\in{\mathcal{A}}_{m}}|\det(a,A_{-i})|$. We use set ${\mathcal{I}}_{m}$ to record the indices of columns where the replacement fails. According to the conditions in line 4, the second for-loop will be repeatedly restarted to ensure that the linear combination of elements in ${\mathcal{B}}_{m}$ can represent all active arms in ${\mathcal{A}}_{m}$ with coefficients in the range of $[-C,C]$. Also note that, as shown in Theorem 4.2, LI-Argmax will be oracle-efficient as it only queries the oracle $\widetilde{\mathcal{O}}\left(d^{3}\right)$ times.
Formally, we have the following result.
Lemma 4.3.
Algorithm 5 Linear-independent argmax (LI-Argmax)
Query oracle to obtain $a^{+}$, the solution of
Query oracle to get $a^{-}$ from (5) by replacing $u$ with $-u$.
4.3 Computational Analysis
Definition 4.4 (Optimization oracle).
We note that the constrained optimization oracles are commonly-used and also crucial for elimination-type approaches e.g., (Bibaut et al., 2020; Li et al., 2022). Although this oracle is slightly powerful than that used for non-elimination based approaches (Dani et al., 2008; Agarwal et al., 2014), linear constrained oracle, in fact, can be implemented efficiently in many cases (e.g., a common assumption that ${\mathcal{A}}$ is a ball (Plevrakis and Hazan, 2020)) via the ellipsoid method. We refer readers to (Bibaut et al., 2020; Plevrakis and Hazan, 2020) for more discussions.
However, the evolution of phases blows up the number of constraints in (10) (e.g., when $m=\Omega(\exp(d))$), which prevents LI-Argmax from directly querying the optimization oracle defined in Definition 4.4, since it requires the number of constraints at most $\mathcal{O}(d)$. To address this issue, one needs to remove the dependence on $m$.
Definition 4.5 ($C$-bounded spanner).
[…]
where $c_{s}$ is the coefficient associated with element $s$.
We enlarge ${\mathcal{A}}_{m-1}$ to ${\mathcal{A}}\cap\text{Span}_{[-C,C]}({\mathcal{B}}_{m-1})$, and one can easily verify that this is true since ${\mathcal{A}}_{m-1}\cap{\mathcal{A}}={\mathcal{A}}_{m-1}$ and Lemma 4.3 implies that ${\mathcal{A}}_{m-1}\subseteq\text{Span}_{[-C,C]}({\mathcal{B}}_{m-1})$.
Lemma 4.6.
[…]
Hence, LI-Argmax can invoke the optimization oracle in Definition 4.4 to solve problems of $\operatorname*{\arg\!\max}_{a\in{\mathcal{A}}_{m-1}}\langle\widehat{\theta}_{m},a\rangle$ and (9) approximately by solving (7) and (5), respectively.
Computational efficiency. We mainly focus on two costly steps in Algorithm 4. The first step is to use Gaussian elimination to obtain the matrix $M_{m}\in\mathbb{R}^{n\times d}$ for some $n\leq d$. This step can be done with at most $\mathcal{O}(d^{3})$ complexity and is implemented at most once for each phase $m$. The other expensive step is to query the optimization oracle to construct a $C$-approximate barycentric spanner. The number of calls to the oracle depends on the number of times that Algorithm 4 invokes LI-Argmax subroutine. The LI-Argmax will be invoked for $d$ times in the first for-loop. For the second for-loop, LI-Argmax will be invoked for $\widetilde{\mathcal{O}}\left(d^{3}\right)$ times. Compared with $\widetilde{\mathcal{O}}\left(d^{2}\right)$ in (Awerbuch and Kleinberg, 2008), the extra $d$ in our complexity comes from the additional restart-condition $i\in{\mathcal{I}}_{m}$ in line 4 of Algorithm 4. As $|{\mathcal{I}}_{m}|\leq d$ and ${\mathcal{I}}_{m}$ is non-increasing when update, this condition will be met for at most $d$ times.
5 Conclusions
@@
In this paper, we propose a new performance measure, the uniform last-iterate (ULI) guarantee, which captures both the instantaneous and cumulative performance of bandit algorithms. To answer the question of whether ULI is (optimally) achievable, we examine three types of algorithms in the finite-arm setting. Specifically, we provide stronger analysis to show that elimination-based algorithms naturally achieve the near-optimal ULI guarantee. We also provide a reduction-based approach to enable any high-probability adversarial algorithm, with a mild condition, to achieve the near-optimal ULI guarantee. We further provide a negative result for optimistic algorithms showing that they cannot be near-optimal with the ULI measure. Finally, in the large arm space setting, we propose an oracle-efficient linear bandit algorithm, equipped with the novel adaptive barycentric spanner technique.
Acknowledgements
YL is supported in part by NSF grant 2221871. LY is supported in part by NSF grant 2221871, and an Amazon Research Grant.
@@ Acknowledgements
$R_{T}$ $=\sum_{t=1}^{T}\Delta_{A_{t}}$
$=\sum_{t=1}^{t_{0}}\Delta_{A_{t}}+\sum_{t=t_{0}+1}^{T}\Delta_{A_{t}}$
@@
where the first inequality holds as we assume all suboptimal gaps are bounded by one, and $\Delta_{A_{t}}\leq F_{\text{ULI}}(\delta,t)\leq G(\delta,t)$ for all $t\geq t_{0}$.
Since $\frac{1}{2}-\alpha\in(0,\frac{1}{2})$, there will be a contradiction to the regret lower bound (Lattimore and Szepesvári, 2020, Corollary 17.3) for the Gaussian MAB setup.
@@
[…]
@@
[…]
@@ B.1 Proof of the first bullet: ULI guarantee implies uniform-PAC bound
$A=\left\{\forall t\in\mathbb{N}:\Delta_{A_{t}}\leq F_{\text{ULI}}(\delta,t)\right\},$ (12)
@@
$B=\left\{\forall\epsilon>0:\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{A_{t}}>\epsilon}\right\}}\leq F_{\text{UPAC}}(\delta,\epsilon)\right\}.$
@@
As a result, the number of times that $\Delta_{A_{t}}>\epsilon$ occurs is at most
$\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{A_{t}}>\epsilon}\right\}}\leq\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{F_{\text{ULI}}(\delta,t)>\epsilon}\right\}}\leq t_{0}\leq\frac{\mathtt{polylog}^{\frac{1}{\kappa}}(t_{0}/\delta)}{\epsilon^{\frac{1}{\kappa}}}.$
@@
$\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{A_{t}}>\epsilon}\right\}}\leq t_{0}\leq\frac{\mathtt{polylog}^{\frac{1}{\kappa}}(t_{0}/\delta)}{\epsilon^{\frac{1}{\kappa}}}\leq\frac{\mathtt{polylog}\left(\delta^{-1}\left(\frac{c_{0}\cdot\mathtt{polylog}(1/\delta)}{\epsilon}\right)^{\frac{1}{\kappa-z}}\right)}{\epsilon^{\frac{1}{\kappa}}}\triangleq F_{\text{UPAC}}(\delta,\epsilon),$ (14)
@@
Case 2: $\epsilon\geq U$. Conditioning on event $A$ and by the definition of $U$, we have $\Delta_{A_{t}}\leq F_{\text{ULI}}(\delta,t)\leq U\leq\epsilon$ for all $t\in\mathbb{N}$, which implies $\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{A_{t}}>\epsilon}\right\}}=0$. Therefore, any positive function works there and we still choose $F_{\text{UPAC}}(\delta,\epsilon)$ the same as above.
@@ B.2 Proof of the second bullet: ULI implies regret bounds
where the last inequality holds by taking the minimum of Eq. (13) and Eq. (14). Therefore, the algorithm will incur no regret, i.e., $\Delta_{A_{t}}=0$ for all $t\geq t_{0}$ For every $T\in\mathbb{N}$, the regret $R_{T}$ is bounded by
$R_{T}$ $=\sum_{t=1}^{T}\Delta_{A_{t}}$
@@
$R_{T}=\sum_{t=1}^{T}\Delta_{A_{t}}\leq\sum_{t=1}^{T}\mathtt{polylog}(t/\delta)\cdot t^{-\kappa}\leq\mathtt{polylog}(T/\delta)\sum_{t=1}^{T}t^{-\kappa}\leq\frac{T^{1-\kappa}\cdot\mathtt{polylog}(\delta^{-1}T)}{1-\kappa}.$

### v2

@@ 4.1 Main Algorithm and Main Results
Theorem 4.2 also shows that, for each phase, the spanner can be constructed by calling the oracle for only polynomial times. Compared to the computational efficiency of the algorithm in (Awerbuch and Kleinberg, 2008), the efficiency of our algorithm is only $d$ times worse than theirs.
5 Achieving Near-Optimal ULI in Tabular Episodic MDPs
Algorithm 3 Tabular Episodic MDPs with ULI guarantee
Input: $\delta\in(0,1)$, set $\Pi_{\texttt{all}}$ containing all deterministic policies, absolute constant $c_{1},c_{2}>0$.
Initialize: $\Pi_{1}=\Pi_{\texttt{all}}$.
for $m=1,2,\ldots$ do
       Set $\delta_{m}=\delta/(2m^{2})$ and duration $T_{m}=\left\lceil c_{1}2^{2m}S^{2}AH^{4}\log^{2}\left(c_{2}2^{2m}S^{2}AH^{4}|\Pi_{\texttt{all}}|/\delta_{m}\right)\right\rceil$. Run Algorithm 4 with input $(\delta_{m},\Pi_{m},T_{m})$ to obtain $\{\widetilde{V}_{m}^{\pi}\}_{\pi\in\Pi_{m}}$. Update active policy set $\Pi_{m+1}=\left\{\pi\in\Pi_{m}:\max_{\pi^{\prime}\in\Pi_{m}}\widetilde{V}_{m}^{\pi^{\prime}}-\widetilde{V}_{m}^{\pi}\leq 2^{-m}\right\}.$
In this section, we propose a novel algorithm that achieves a near-optimal ULI guarantee in tabular episodic MDPs setup. The algorithm is formally presented in Algorithm 3.
High-level idea. Algorithm 3 conducts policy elimination over a policy set $\Pi_{\texttt{all}}$ which enumerates all deterministic policies. The key challenge here is to acquire accurate estimations of value functions uniformly for all deterministic policies. Note that naïvely playing all deterministic policies will incur linear dependence on $|\Pi_{\texttt{all}}|=A^{SH}$ in $F_{\text{ULI}}$, which is exponentially large. Hence, the algorithm invokes subroutine Algorithm 4 to exhaustively explore the environment, which ensures accurate estimations of the transition model and only suffers logarithmic dependence on $|\Pi_{\texttt{all}}|$. Once the transition model can be well-approximated, the algorithm constructs an accurate estimation of the value function for every policy and then decides which policies should be eliminated.
More concretely, Algorithm 3 first accepts a set of all deterministic policies and then it proceeds in phases $m=1,2,\ldots$. In each phase $m$, subroutine Algorithm 4 is invoked to learn the transition model. Specifically, the subroutine inherits the structure of UCB-VI (Azar et al., 2017), but more importantly the algorithm pretends to be agnostic to the reward function and uses uncertainty-driven reward functions $\{b_{t,h}(s,a)/H\}_{t,h}$ where
$b_{t,h}(s,a)=H\sqrt{\frac{2S\log\iota}{\max\{1,N_{t,h}(s,a)\}}}+\frac{2HS\log\iota}{3\max\{1,N_{t,h}(s,a)\}}\text{ where }\iota=\frac{10SAH|\Pi_{\texttt{all}}|T}{\delta},$ (5)
where $N_{t,h}(s,a)$ is the number of times of visiting $(s,a)$ at stage $h$ up to episode $t$. Note that Eq. (5) captures the uncertainty of visitation a state-action pair, i.e., more visitations of $(s,a)$, larger $N_{t,h}(s,a)$, and less uncertainty. This modification, inspired by (Wang et al., 2020) forces the algorithm to aggressively explore the environment, in the sense that the algorithm prefers to play a policy that maximizes the uncertainty.
Algorithm 4 Uniform estimation for value functions
Input: confidence $\delta\in(0,1)$, policy set $\Pi\subseteq\Pi_{\texttt{all}}$, duration $T$.
Initialize: randomly pick a policy $\pi_{1}\in\Pi$, $N_{1,h}(s,a)=N_{1,h}(s,a,s^{\prime})=0$ for all $(h,s,a,s^{\prime})$.
for $t=1,\ldots,T$ do
       Observe initial state $s_{t,1}\sim\mu$. for $h=1,\ldots,H$ do
             Take action $a_{t,h}=\pi_{t,h}(s_{t,h})$ and observe $s_{t,h+1}\sim\mathbb{P}_{h}(\cdot|s_{t,h},a_{t,h})$. Increase counters $N_{t,h}(s_{t,h},a_{t,h},s_{t,h+1})\overset{+}{\leftarrow}1$ and $N_{t,h}(s_{t,h},a_{t,h})\overset{+}{\leftarrow}1$.
      Update estimates ^$\widehat{\mathbb{P}}_{t,h}(s^{\prime}|s,a)=\frac{N_{t,h}(s,a,s^{\prime})}{\max\{1,N_{t,h}(s,a)\}}$ for all $(s,a,s^{\prime})\in{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}$. Update bonus function $b_{t}=\left\{b_{t,h}\right\}_{h\in[H]}$ where $b_{t,h}(\cdot,\cdot)$ is updated according to Eq. (5). Get $\{\widehat{V}^{\pi}_{t,1}(s_{1})\}_{\pi\in\Pi}$ by invoking Algorithm 5 with input $(\Pi,b_{t}/H,\{\widehat{\mathbb{P}}_{T,h}\}_{h\in[H]},b_{t},s_{t,1})$. Update policy $\pi_{t+1}=\operatorname*{\arg\!\max}_{\pi\in\Pi}\widehat{V}^{\pi}_{t,1}(s_{1})$.
for $t=1,\ldots,T$ do
       Get $\{\widehat{V}^{\pi}_{T,1}(s_{t,1})\}_{\pi\in\Pi}$ by invoking Algorithm 5 with input $(\Pi,r,\{\widehat{\mathbb{P}}_{T,h}\}_{h\in[H]},0,s_{t,1})$.
Output: $\{\widetilde{V}^{\pi}\}_{\pi\in\Pi}$ where $\widetilde{V}^{\pi}=\frac{1}{T}\sum_{t=1}^{T}\widehat{V}^{\pi}_{T,1}(s_{t,1})$.
Algorithm 5 Construct estimated value function
Input: policy set $\Pi\subseteq\Pi_{\texttt{all}}$, reward function $r$, transition $\widehat{\mathbb{P}}$, bonus function $b$, initial state $s_{1}$.
for $\pi\in\Pi$ do
       $\widehat{Q}^{\pi}_{H+1}(\cdot,\cdot)=0$ and $\widehat{V}^{\pi}_{H+1}(\cdot)=0$. for $h=H,H-1,\ldots,1$ do
             $\widehat{Q}^{\pi}_{h}(\cdot,\cdot)=\min\left\{\left[\widehat{\mathbb{P}}_{h}\widehat{V}^{\pi}_{h+1}\right](\cdot,\cdot)+r_{h}(\cdot,\cdot)+b_{h}(\cdot,\cdot),H\right\}$ and $\widehat{V}^{\pi}_{h}(\cdot)=\widehat{Q}^{\pi}_{h}(\cdot,\pi(\cdot))$.
Output: $\{\widehat{V}^{\pi}_{1}(s_{1})\}_{\pi\in\Pi}$.
The following theorem shows that our algorithm achieves the ULI guarantee.
Theorem 5.1.
For any fixed $\delta\in(0,1)$, Algorithm 3 achieves the ULI guarantee with function $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{S^{3}AH^{5}}\log\left(tSAH/\delta\right)\big{)}$.
Theorem 5.1 shows near-optimality of Algorithm 3 w.r.t. episode $t$. Unfortunately, the regret bound implied by our ULI result incurs suboptimal dependence on $S,H$, and the logarithmic term, and our algorithm is not computational efficient. We leave these improvements for future work.
6 Conclusions and Future Work
@@
In this paper, we propose a new metric, the uniform last-iterate (ULI) guarantee, which captures both instantaneous and cumulative performance of sequential decision-making algorithms. To answer whether ULI is (optimally) achievable, we first examine three types of bandit algorithms in the finite-arm setting. Specifically, we provide stronger analysis to show that elimination-based algorithms naturally achieve near-optimal ULI guarantees. We also provide a reduction-based approach to enable any high-probability adversarial algorithms, with a mild condition, to achieve near-optimal ULI guarantees. We further provide a negative result for optimistic bandit algorithms showing that they cannot achieve near-optimal ULI guarantee. Furthermore, in the large arm space setting, we propose an oracle-efficient linear bandit algorithm, equipped with the novel adaptive barycentric spanner technique. Finally, we propose a new algorithm, which adapts uncertainty-driven reward functions into policy elimination to achieve a ULI guarantee in tabular episodic MDPs.
Some natural directions are to improve our current results, including proposing a ULI algorithm for linear bandits with infinite arms and proposing a computationally efficient algorithm, possibly based on the action-elimination approach, for tabular MDPs. We also summarize other interesting directions as follows.
  • Design (computationally efficient) algorithms for MDPs with linear function approximation. The main challenge is to bypass any dependence on the number of states which is possibly infinite. Thus, generalizing our RL algorithm, which enumerates all deterministic policies, to linear MDPs does not work.
  • Design (computationally efficient) algorithms for Episodic MDPs with only logarithmic dependence on $H$ (a.k.a. horizon-free). In our attempts, we use the doubling trick on $\epsilon$ in existing $(\delta,\epsilon)$-PAC horizon-free RL algorithms. In this case, we run the algorithm in phases with $\epsilon,\epsilon/2,\epsilon/4,\ldots$. The main difficulty is to leverage the information learned from the previous phase to guide the algorithm to play an improved policy at the next phase as required by ULI. Thus, we conjecture that there might exist fundamental barriers to simultaneously achieve ULI guarantee and a logarithmic dependence on $H$.
  • It could be also interesting to investigate some empirical issues when deploying ULI algorithms in high-stakes fields. Typically, optimistic algorithms initially incur a lower regret than ULI algorithms, but their regret will exceed those of ULI algorithms at a certain juncture. At this juncture, ULI algorithms have eliminated all suboptimal arms, whereas optimistic algorithms continue exploring as time evolves. Hence, identifying this turning point could be beneficial for deploying ULI algorithms in high-stakes domains.
@@
$R_{T}$ $=\sum_{t=1}^{T}\Delta_{t}$
$=\sum_{t=1}^{t_{0}}\Delta_{t}+\sum_{t=t_{0}+1}^{T}\Delta_{t}$
@@
where the first inequality holds as we assume all suboptimal gaps are bounded by one, and $\Delta_{t}\leq F_{\text{ULI}}(\delta,t)$ for all $t\geq t_{0}$.
Since $\frac{1}{2}-\alpha\in(0,\frac{1}{2})$, there will be a contradiction to the regret lower bound [Lattimore and Szepesvári, 2020, Corollary 17.3] for the Gaussian MAB setup.
@@
Suppose an algorithm achieves the ULI guarantee for some $\delta$ with function $F_{\text{ULI}}(\delta,t)=\mathtt{polylog}(t/\delta)\cdot t^{-\kappa}$ where $\kappa\in(0,1)$. Then, we have,
@@
    $F_{\text{UPAC}}(\delta,\epsilon)=\mathcal{O}\left(\frac{\mathtt{polylog}(\delta^{-1}\epsilon^{-1})}{\epsilon^{1/\kappa}}\right).$
@@ B.1 Proof of the first bullet: ULI guarantee implies uniform-PAC bound
$A=\left\{\forall t\in\mathbb{N}:\Delta_{t}\leq F_{\text{ULI}}(\delta,t)\right\},$ (6)
@@
$B=\left\{\forall\epsilon>0:\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{t}>\epsilon}\right\}}\leq F_{\text{UPAC}}(\delta,\epsilon)\right\}.$
@@
As a result, the number of times that $\Delta_{t}>\epsilon$ occurs is at most
$\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{t}>\epsilon}\right\}}\leq\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{F_{\text{ULI}}(\delta,t)>\epsilon}\right\}}\leq t_{0}\leq\frac{\mathtt{polylog}^{\frac{1}{\kappa}}(t_{0}/\delta)}{\epsilon^{\frac{1}{\kappa}}}.$
@@
$\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{t}>\epsilon}\right\}}\leq t_{0}\leq\frac{\mathtt{polylog}^{\frac{1}{\kappa}}(t_{0}/\delta)}{\epsilon^{\frac{1}{\kappa}}}\leq\frac{\mathtt{polylog}\left(\delta^{-1}\left(\frac{c_{0}\cdot\mathtt{polylog}(1/\delta)}{\epsilon}\right)^{\frac{1}{\kappa-z}}\right)}{\epsilon^{\frac{1}{\kappa}}}\triangleq F_{\text{UPAC}}(\delta,\epsilon),$ (8)
@@
Case 2: $\epsilon\geq U$. Conditioning on event $A$ and by the definition of $U$, we have $\Delta_{t}\leq F_{\text{ULI}}(\delta,t)\leq U\leq\epsilon$ for all $t\in\mathbb{N}$, which implies $\sum\limits_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{t}>\epsilon}\right\}}=0$. Therefore, any positive function works there and we still choose $F_{\text{UPAC}}(\delta,\epsilon)$ the same as above.
@@ B.2 Proof of the second bullet: ULI implies regret bounds
where the last inequality holds by taking the minimum of Eq. (7) and Eq. (8). Therefore, the algorithm will incur no regret, i.e., $\Delta_{t}=0$ for all $t\geq t_{0}$ For every $T\in\mathbb{N}$, the regret $R_{T}$ is bounded by
$R_{T}$ $=\sum_{t=1}^{T}\Delta_{t}$
@@
$R_{T}=\sum_{t=1}^{T}\Delta_{t}\leq\sum_{t=1}^{T}\mathtt{polylog}(t/\delta)\cdot t^{-\kappa}\leq\mathtt{polylog}(T/\delta)\sum_{t=1}^{T}t^{-\kappa}\leq\frac{T^{1-\kappa}\cdot\mathtt{polylog}(\delta^{-1}T)}{1-\kappa}.$